IF A BAG CONTAINS RED, BLUE AND ORANGE MARBLES AND THE PROBABILITY OF GETTING A RED MARBLE IS 0.4 AND THE PROBABILITY OF GETTING A RED OR BLUE IS 0.9 WHAT IS THE PROBABILITY OF GETTING A RED OR ORANGE
Answer by nyc_function(2741)

Let R = the probability of getting a red ball, B the probability of getting a blue, and O the probability of getting an orange one.
Because you have to get something, the sum of the probabilities is R + B + O = 1.0
You are given
R = 0.4
R + B = 0.9
These two equations together give that B = 0.5
Now that you know R and B, O must be equal to 0.1 for the sum of the probabilities to be 1.0
So, then R + O = 0.5 <--answer.